# Constitutional Consensus

### v4

@@
Abstract
Consider people with smartphones operating without external authorities or global resources other than the network itself. In this setting, high-end applications supporting sovereign democratic digital communities, community banks, and digital cooperatives require consensus executed by community members, which must be reconfigurable to support community dynamics.
The Constitutional Consensus protocol aims to address this need by introducing constitutional self-governance to consensus: participants dynamically amend the participant set, supermajority threshold, and timeout parameter through the consensus protocol itself. We achieve this by enhancing a DAG-based protocol (like Cordial Miners) with participant-controlled reconfiguration, while also supporting both high- and low-throughput operation (like Morpheus), remaining quiescent when idle. This three-way synthesis uniquely combines: (1) constitutional amendments for self-governance, (2) a cryptographic DAG structure for simplicity, parallelism, and throughput, and (3) both high- and low-throughput operation. The protocol achieves consensus in $3\delta$, maintains $O(n)$ amortized communication complexity during high throughput, and seamlessly transitions between modes. The basic protocol (without constitutional amendments) realizes these features in 25 lines of pseudocode, making it one of the most concise consensus protocols for eventual synchrony.
keywords:
@@ Abstract
Consensus Protocols, Grassroots Platforms, Blocklace, Digital Social Contracts, Democratic Digital Communities
1 Introduction
Participant-Controlled Consensus. Consider people who wish to form self-governing digital communities without relying on servers or external authorities. These participants—whether members of community banks, digital cooperatives, or democratic assemblies—need consensus protocols they can execute on their smartphones while maintaining complete control over membership and governance rules. Unlike traditional reconfiguration where external authorities determine changes [40, 2], these communities require constitutional self-governance: participants must collectively control who joins, how decisions are made, and how the protocol operates, all through the consensus protocol itself. This paper presents Constitutional Consensus, a protocol that enables such self-governance by allowing participants to dynamically amend the participant set $P$, supermajority threshold $\sigma$, and timeout parameter $\Delta$ through democratic decisions implemented within the protocol.
Digital Social Contracts. These self-governing communities implement digital social contracts [9]—deterministic state machines that encode agreements among people who know each other. Like smart contracts, they provide computational guarantees about contract execution. However, while smart contracts [10] operate among anonymous accounts and run on consensus protocols executed by third-party servers, social contracts are among identified people and run on consensus protocols executed by the participants’ own devices. Existing consensus-based smart contracts, notably in Decentralized Finance (DeFi) platforms [10, 7, 25], have driven development of high-performance consensus protocols [43, 17, 18, 4, 16], but these operate among pseudonymous accounts on global networks rather than identified people in local communities. This fundamental difference enables democratic rather than plutocratic governance, as shown in Table 1.
Table 1: Comparing Smart Contracts and Consensus-Based Digital Social Contracts
Smart Contract [10] Digital Social Contract [9]
Among: Accounts People
Executed by: Third-party servers The people’s smartphones
In return for: Gas Naught
Applications: Decentralized Finance (DeFi) [31, 15],
Decentralized Autonomous Organizations (DAOs) [12, 11],
Non-Fungible Tokens (NFTs) [41]
Grassroots democratic digital communities [28, 33] and their federation [38, 39],
community banks and digital cooperatives [37]
Amended via: Self-modifying code Constitutional amendments
Governance: Token-weighted voting (plutocracy) One person, one vote (democracy)
Amendment Authority: External governance tokens or DAOs Participants themselves via supermajority
Parameter Control: Fixed in code or controlled by token holders Democratically adjustable by participants
We note that while smart contracts are necessarily consensus-based, digital social contracts are not. Some grassroots platforms, notably grassroots social networks [35] and grassroots cryptocurrencies [37, 22], do not require consensus on the total ordering of the acts of their participants: Their safety depends only on the agents obeying the social contract [38]. This can be ensured by mechanisms weaker then consensus, e.g. mutual attestations among agents operating within Trusted Execution Environments (TEEs) [30]. However, applications requiring collective agreement—such as community banks with multiple signatories or democratic governance with voting—do require consensus to ensure all participants agree on the order of transactions and decisions.
Technical Requirements. Communities implementing digital social contracts face unique technical challenges. Small communities may have transaction rates much lower than network latency, requiring a protocol that remains quiescent—sending no messages when idle. Yet the same protocol must efficiently handle high-throughput periods when the community is larger and/or more active. During low throughput, transactions arrive one-at-a-time with no conflicts to resolve, so ideally any participant should be able to finalize their transaction without designated leaders. During high throughput, the protocol should match the performance of modern consensus protocols.
Specifically, an effective protocol for self-governing communities requires:
  1. 1.
    Quiescent operation: No activity when idle, spontaneous leader emergence for isolated transactions
  2. 2.
    Dual-mode efficiency: Seamless transition between low and high throughput modes
  3. 3.
    Responsiveness: Network-speed operation unaffected by conservative timeout estimates
  4. 4.
    DAG-based structure: Leveraging blocklace [3] parallelism and integration
  5. 5.
    Constitutional amendments: Participant-controlled governance of protocol parameters
Technical Foundation. Constitutional Consensus achieves these requirements through a three-way synthesis. From Morpheus [23], it inherits dual-mode operation with spontaneous leaders during low throughput and formal leadership during high throughput. From Cordial Miners [18], it adopts the blocklace structure for parallelism and efficiency. To these, it adds constitutional amendment mechanisms enabling true self-governance. The protocol operates in waves of three rounds. During low throughput, any participant with a transaction can spontaneously lead a wave, achieving finality in $3\delta$. During high throughput, formal leaders are selected round-robin, maintaining $O(n)$ amortized communication complexity. The basic protocol (without amendments) requires only 25 lines of pseudocode, making it one of the most concise consensus protocols for eventual synchrony.
Constitutional Self-Governance. The key innovation lies in how participants control their own governance. Unlike smart contracts that rely on external governance tokens or DAOs [12], constitutional amendments are decided by the participants themselves through supermajority vote. While Lamport’s original Paxos envisioned a self-governing parliament [21], subsequent reconfiguration protocols [42, 2, 40, 20] assume external authorities determine membership changes. Constitutional Consensus returns to Paxos’s original vision of true self-governance. Unlike Proof-of-Stake systems where validators must stake capital [8], participation requires only the democratic approval of existing members. This creates a fundamental architectural difference: while blockchain systems typically support only a single global instance whose security depends on token value, Constitutional Consensus enables unlimited independent communities to self-govern simultaneously.
Connection to Grassroots Architecture. This capability for multiple independent instances operating without central coordination or global resources makes Constitutional Consensus inherently grassroots [38]. A community in Bangladesh, a cooperative in Brazil, and a federation in Berlin can each run their own instance, potentially with overlapping members, without any coordination or competition for resources. This architecture—multiple local instances rather than a single global instance—enables the broader grassroots vision of egalitarian digital platforms that operate without central authorities [36, 34]. Within this grassroots architecture, Constitutional Consensus provides the consensus layer for applications requiring collective agreement, complementing consensus-free grassroots platforms like social networks [35] and cryptocurrencies [37, 22]. Together, these protocols enable communities to build sovereign democratic digital infrastructure from the ground up, controlled entirely by their members rather than external authorities or token holders. This distinguishes Constitutional Consensus from global platforms—whether centralized like Facebook or decentralized like Bitcoin [26] and IPFS [5]—that can support only a single instance.
Example: Grassroots Communities, Banks, and Federations. As a concrete scenario for grassroots platforms, consider a small town whose residents wish to form a digital community. They begin by deploying independent instances of consensus-free grassroots platforms on their smartphones—creating their own grassroots social network [35] and trading via self-minted grassroots cryptocurrencies [37, 22]. Crucially, they do this without coordination with any global network or central authority. When their needs evolve to require collective agreement—community banks with multiple signatories, constitutional governance, or federation with neighbouring towns—they deploy their own Constitutional Consensus instance. This instance operates independently yet can interoperate with other communities’ instances, embodying the grassroots principle of multiple local instances rather than a single global system. Through Grassroots Federation [39], these autonomous instances can scale democratically from local to potentially global governance. Small communities ($\approx 100$ members) govern themselves directly; larger ones use sortition-selected assemblies [14, 39]. Each community’s governance runs as a digital social contract on their own Constitutional Consensus instance, executed by members on their smartphones—no external servers or authorities required.
Model. As deterministic fault-tolerant consensus is not solvable in asynchronous systems [13], we assume an eventual synchrony model, where communication delays are at most $\delta$ after some Global Stabilization Time (GST) that is unknown to the agents. We consider Byzantine failures, as some malicious actors may infiltrate the network, but we assume a bounded ratio of Byzantine agents (to be specified later), since, by and large, the agents trust each other.
Paper structure. Section 2 provides the model and problem definitions. Section 3 presents a single-epoch Constitutional Consensus protocol, with only an initial constitution and no constitutional amendments. Section 4 proves the safety and liveness of single-epoch Constitutional Consensus. Section 5 adds support for constitutional amendments – for each of the three parameters $P,\sigma$, and $\Delta$, we present a democratic process for amending it. We then present the protocol extension needed to realize constitutional amendments and prove its safety and liveness. Section 6 proves that Constitutional Consensus is grassroots. Section 7 concludes.
2 Model and Problem
@@
Model. We assume a set of agents $\Pi$, each endowed with a unique key-pair and identified by its public key $p\in\Pi$. While $\Pi$ is potentially-infinite, we refer only to finite subsets $P\subset\Pi$ of it. A correct agent follows the protocol. An agent that does not follow the protocol is faulty or Byzantine.
@@
We assume an eventual synchrony (aka partial synchrony) model in which every message sent between correct agents is eventually received. Message latency is unbounded, but in every run there is a time called the Global Stabilization Time (GST), which is unknown to the agents, after which all messages arrive within a least upper bound $\delta$ time for some $\delta>0$.
Note that the abstraction of reliable messaging does not mean that agents must communicate via point-to-point TCP connections. Rather, it abstracts away a dissemination layer that can be implemented using acks/nacks, gossip, network coding, etc., and is orthogonal to the consensus protocol in this paper.
@@
A protocol proceeds through a sequence of epochs, each commences with a new constitution (defined next), referred to as the prevailing constitution during the epoch, and ends when the constitution is amended, namely replaced with a new constitution. The constitution defines the set of participants in the epoch, as well as key protocol parameters:
Definition 2.1 (Constitution, $\sigma$-Supermajority).
A constitution is a triple $(P,\sigma,\Delta)$ where:
  1. 1.
    $P\subset\Pi$ is a set of agents, with $n:=|P|$,
  2. 2.
@@
    $\frac{1}{2}\leq\sigma<1$ is a fraction that specifies a supermajority among the agents $P$,
  3. 3.
@@
    $\Delta>0$ is a presumed upper-bound on message delay after GST among $P$.
@@
We may omit $\Delta$ when its value is immaterial (e.g., when discussing protocol safety) and refer to the constitution $(P,\sigma)$. A $\sigma$-supermajority among $P$ is a fraction $Q\subseteq P$ such that $|Q|>\sigma n$.
Supermajorities and faulty agents. Typically, a permissioned consensus protocol assumes that there are at most $f<n$ faulty agents among $P$, and with this assumption we define $\sigma:=\frac{n+f}{2n}$. Equivalently, $\sigma$ is chosen with the assumption that there are at most $f=(2\sigma-1)n$ faulty agents in $P$. For example, for $f=0$, $\sigma=\frac{1}{2}$; for $f=\frac{1}{2}n$, $\sigma=\frac{3}{4}$; and if everyone but one agent is faulty then $\sigma=\frac{2n-1}{2n}=1-\frac{1}{2n}$, namely unanimity is required. For $n=3f+1$ (the standard assumption), $\sigma=\frac{n+\frac{n-1}{3}}{2n}=\frac{2}{3}-\frac{1}{6n}$, namely a supermajority greater-or-equal to $\frac{2}{3}n$ is required.
Typically, permissioned consensus protocols wait to hear from “at least nfn-f agents among PP”, which in the standard case of $n=3f+1$ (but not otherwise) happens to be equivalent to “a $\sigma$-supermajority among PP”.
@@
Constitution safety and liveness. A constitution thus defines the protocol’s assumptions about the environment: $P$ is the set of agents running the protocol; there is a correct $\sigma$-supermajority in $P$ and every $\sigma$-supermajority in $P$ includes a majority of correct agents, hence the intersection of any two $\sigma$-supermajorities contains a correct agent; $\Delta\geq\delta$ is an upper bound on the communication time after GST.
Definition 2.2 (Safe and Live Constitution).
A constitution $(P,\sigma,\Delta)$ is safe if $P$ includes at most $f=(2\sigma-1)n$ faulty agents, and live if after GST any message sent among correct agents in $P$ arrives within $\Delta$.
@@
Constitutional Consensus problem definition. A constitutional consensus protocol may have multiple instances. Each instance commences with an initial constitution $(P,\sigma,\Delta)$ that defines its participating agents $P$ and the key protocol parameters $\sigma$ and $\Delta$. As in standard consensus, participants spontaneously issue transactions (these are the inputs to the protocol) and every agent outputs a sequence of transactions.
@@
In addition, constitutional consensus assumes an that participants engage in an unspecified democratic decision process that may result in a decision $d$ to amend the prevailing constitution. The discussion of such democratic processes is deferred to Section 5. For the purpose of the constitutional consensus protocol, all is needed is to assume: (i) the existence of such a process that may result in constitutional amendment decisions, which become known to the participants once made; (ii) that participants can verify a constitutional amendment decisions $d$ to be valid according to the rules specified in Section 5. A valid constitutional amendment decision $d$ defines a new constitution $(P^{\prime},\sigma^{\prime},\Delta^{\prime})$, in which the participants $P^{\prime}$ and/or the parameters $\sigma^{\prime}$ and $\Delta^{\prime}$ are the result of amending the prevailing constitution.
@@
In the following, constitution refers to the new constitution of a valid constitutional amendment decision, and an initial constitution is the new constitution of a valid amendment of a constitution with an empty population. Valid amendments are provided to the consensus protocol as inputs by participants that know of them. Thus, the output sequence of a constitutional consensus protocol includes both transactions and valid constitutional amendment decisions, as follows:
@@
The constitutional output sequences produced by correct agents executing a constitutional consensus protocol instance must satisfy the following safety and liveness conditions:
@@
In any run of an instance of constitutional consensus in which all valid constitutional amendment decisions made are safe and live, there exists a constitutional sequence $T=t_{1},t_{2},\dots$ for which the following safety and liveness properties hold:
  1. 1.
    Safety:
    Consistency
@@
    Validity
    If a correct agent $p$ outputs a transaction $t$, then $p$ is a participant in $t$’s epoch in $T$.
  2. 2.
    Liveness:
    Epoch Liveness
    If a correct agent $p$ is a participant in an epoch $d$ in $T$, then $p$’s output sequence eventually includes all the transactions in $d$’s epoch sequence. Furthermore, if $d$ is the last epoch in $T$, then every transaction input by a correct participant in the epoch $d$ is included in $T$.
    Constitution Amendment Liveness
@@
Note that if there is no decision to amend the initial constitution then the protocol run has a single epoch with a fixed set of participants. In this case, the definition above coincides with the standard consensus correctness definition.

### v5

@@
Abstract.
Permissionless-consensus-based Decentralised Autonomous Organisations (DAOs) are the prevailing paradigm for participant-governed digital organisations. As participants have verified resources but no trusted identities, this ecosystem is necessarily plutocratic (one coin – one vote). Here we offer, for the first time, a democratic (one person – one vote) paradigm for the governance of digital communities and organisations, based on permissioned consensus and egalitarian decision processes.
In line with Lamport’s vision of consensus as a self-governing parliament, in the democratic paradigm a constitution specifies both a decision making protocol as well as a consensus protocol, combined to let participants amend the constitution through constitutionally-valid decisions that are ratified by consensus. To meaningfully instantiate this paradigm we integrate the disciplines of distributed computing and computational social choice, with the goal of providing a practical and efficient smartphone-based solution for the democratic self-governance of grassroots sovereign digital communities and organisations.
The resulting Constitutional Consensus protocol employs (1) state-of-the-art Sybil-resilient democratic decision processes for amending the set of participants, supermajority threshold, and timeout; and (2) a novel Byzantine-fault tolerant consensus protocol that is DAG-based (following Cordial Miners) thus eschewing reliable broadcast, with dual-mode operation (following Morpheus) that is quiescent when idle, has spontaneous leaders for isolated transactions, and formal round-robin leadership during high throughput. Constitutional Consensus operates in epochs, each governed by the prevailing constitution, with the next epoch commencing via the ratification of a constitutionally-valid constitutional-amendment decision. The protocol achieves consensus in $3\delta$ with $O(n)$ amortized communication complexity during high throughput. The single-epoch protocol is quite simple, requiring only 25 lines of pseudocode; the multi-epoch protocol is grassroots, as each member can initiate and participate in multiple independent instances that can form, operate and grow concurrently without coordination or dependence on global resources other than the net.
@@ Abstract.
Consensus, Paxos, constitutional amendment, digital social contracts, democratic digital communities, grassroots
$c$s: Computer systems organization Peer-to-peer architecturesccs: Software and its engineering Distributed systems organizing principles
1. Introduction
Participant-governance of digital organisations. Permissionless-consensus-based Decentralised Autonomous Organisations (DAOs) (Ethereum, ndao; Ding et al., 2023; Hassan and De Filippi, 2021; Talmon, 2023) are the prevailing paradigm for realising participant-governed digital organisations. As participants have verified resources (e.g., staked tokens) but no trusted identities, this ecosystem is necessarily plutocratic: voting power is proportional to token holdings (one coin – one vote). Here we offer, for the first time, a democratic (one person – one vote) paradigm for the governance of digital communities and organisations, based on permissioned consensus and egalitarian decision processes.
Constitutional self-governance. Reconfiguration—changing the set of participants in a consensus protocol—has been studied extensively (Lamport, 1998, 2001; Lamport et al., 2009; Aguilera et al., 2011; Spiegelman et al., 2017; Ongaro and Ousterhout, 2014). This literature addresses the mechanism of how membership changes execute safely, but not who decides on them: the decision authority is either left unspecified or delegated to an external entity. Moreover, prior work applies reconfiguration only to membership, leaving the rest of the protocol fixed. [^1] Tezos (Goodman, 2014) has no notion of membership but can self-modify the protocol via plutocratic governance. In line with Lamport’s vision of consensus as a self-governing parliament (Lamport, 1998), we propose a democratic paradigm in which a constitution specifies both a decision-making protocol and a consensus protocol, combined to let participants amend the constitution through constitutionally-valid decisions that are ratified by consensus. To meaningfully instantiate this paradigm we integrate the disciplines of distributed computing (Lynch, 1996) and computational social choice (Brandt et al., 2016), with the goal of providing a practical and efficient smartphone-based solution for the democratic self-governance of grassroots (Shapiro, 2023a) sovereign digital communities and organisations. The resulting Constitutional Consensus protocol presented here employs (1) state-of-the-art Sybil-resilient (Shahaf et al., 2020) democratic decision processes for amending the set of participants (Bulteau et al., 2021), supermajority threshold (Abramowitz et al., 2021a), and timeout (Shahaf et al., 2019); and (2) a novel Byzantine-fault tolerant consensus protocol that is DAG-based (following Cordial Miners (Keidar et al., 2023)) thus eschewing reliable broadcast, with dual-mode operation (following Morpheus (Lewis-Pye and Shapiro, 2025)) that is quiescent when idle, has spontaneous leaders for isolated transactions, and formal round-robin leadership during high throughput.
Digital social contracts. A digital social contract (Cardelli et al., 2020) is a voluntary agreement among a set of people, specified, fulfilled, and enforced digitally. Many digital social contracts — including social graphs (Shapiro, 2026), social networks (Shapiro, 2023b), and digital currencies (Shapiro, 2024; Lewis-Pye et al., 2023)—require no consensus at all. Those that do require consensus among the contract’s participants—notably democratic governance and federation of sovereign communities (Shapiro and Talmon, 2025)—may execute the protocol on the participants’ own devices. Constitutional Consensus is designed for such social contracts. Unlike smart contracts (De Filippi et al., 2021), whose execution is enforced by global consensus among third-party validators on a blockchain, digital social contracts are among people with mutually-known identities and are enforced by the participants themselves. This distinction—summarised in Table 1 in Appendix J—has a direct architectural consequence: governance is democratic rather than plutocratic.
Technical approach. Constitutional Consensus operates in epochs, each governed by the prevailing constitution, with the next epoch commencing via the ratification of a constitutionally-valid constitutional-amendment decision. The single-epoch protocol is quite simple, requiring only 25 lines of pseudocode. After GST, a leader block is finalised within $3\delta$, and the amortized communication complexity per transaction is $O(n)$ during high throughput.
Grassroots. Because Constitutional Consensus requires no external authority, global ledger, or staking globally-controlled capital, the multi-epoch protocol is grassroots (Shapiro, 2023a, 2026): each member can initiate and participate in multiple independent instances that can form, operate and grow concurrently without coordination or dependence on global resources other than the net. Any group of people can bootstrap their own instance on their smartphones. This is in contrast with Proof-of-Stake protocols like Ethereum 2.0 (Buterin et al., 2020), where security depends on the staked token having real-world value, creating a winner-take-all dynamic that precludes independent instances.
Paper structure. Section 2 provides the model and problem definitions. Section 3 presents single-epoch Constitutional Consensus, with no constitutional amendments. Section 4 presents Sybil-resilient democratic decision processes for amending the set of participants, supermajority threshold, and timeout. Section 5 presents the formal amendment mechanism and epoch transitions. Section 6 summarises related work, with a full survey in Appendix J. Section 7 concludes. Correctness proofs and other technicalities are deferred to appendices.
2. Model and Problem
@@
We assume a set of agents $\Pi$, each has a unique key-pair and identified by its public key $p\in\Pi$. While $\Pi$ is potentially-infinite, we refer only to finite subsets $P\subset\Pi$ of it. A correct agent follows the protocol. An agent that does not follow the protocol is faulty or Byzantine.
@@
We assume an eventual synchrony (partial synchrony) model in which every message sent between correct agents is eventually received (this abstracts a dissemination layer that can use acks, gossip, or network coding, and is orthogonal to the consensus protocol). Message latency is unbounded, but in every run there is a time called the Global Stabilization Time (GST), which is unknown to the agents, after which all messages arrive within a least upper bound $\delta$ time for some $\delta>0$.
In general, the constitution specifies the entire set of protocols employed—decision making processes as well as the consensus protocol for ratifying them. In our case, these consist of known democratic decision processes recalled in Section 4 and the novel Constitutional Consensus protocol, which is the main subject of this paper. In this work the constitution assumes both of these protocols to be fixed except for three amendable components, the participant set, the supermajority threshold, and the timeout parameter.
@@
A protocol proceeds through a sequence of epochs, each commences with a new constitution, referred to as the prevailing constitution during the epoch, and ends when the constitution is amended, namely replaced with a new constitution:
Definition 2.1 (Amendable component of the Constitution).
The amendable component of the constitution is the triple $(P,\sigma,\Delta)$ where:
  1. (1)
    $P\subset\Pi$ is a set of agents, referred to as its participants, with $n:=|P|$,
  2. (2)
@@
    $\frac{1}{2}\leq\sigma<1$ is a fraction that specifies a supermajority among the participants,
  3. (3)
@@
    $\Delta>0$ is a presumed upper-bound on message delay after GST among the participants.
@@
We may omit $\Delta$ when discussing protocol safety and refer to the constitution $(P,\sigma)$. A $\sigma$-supermajority among the participants is a fraction $Q\subseteq P$ such that $|Q|>\sigma n$.
In the following we may use the constitution as a shorthand, e.g. “the prevailing constitution” instead of “the amendable component of the prevailing constitution”.
Supermajorities, Sybils and Byzantines. Our protocols are defined using a supermajority that aims to overcome participants that are (1) Byzantine (arbitrarily-behaving) (Lamport et al., 1982), which hamper consensus and (2) Sybil (with fake or duplicate identity) (Shahaf et al., 2020), which hamper equality. Assuming there are at most $f<n$ such faulty participants, we define $\sigma:=\frac{n+f}{2n}$. Equivalently, $\sigma$ is chosen with the assumption that there are at most $f=(2\sigma-1)n$ faulty participants. Both protocols ensure safety with any $f<n$ and liveness if $f<\frac{1}{3}n$. Typically, permissioned consensus protocols wait to hear from “at least nfn-f agents among PP”, which in the standard case of $n=3f+1$ (but not otherwise) happens to be equivalent to “a $\sigma$-supermajority among PP”. While Byzantines and Sybils are related (e.g. Byzantines may aim to create and control multiple Sybils to amplify their ability to cause harm) they are not the same, and a realisation of Constitutional Consensus may choose to incorporate a tighter supermajority for each, say $\sigma_{B}$ for consensus and $\sigma_{S}$ for democratic decisions; for simplicity, we employ here one supermajority $\sigma$ for both.
@@
Constitution safety and liveness. The (amendable component of the) constitution thus defines the protocol’s assumptions about the environment: A constitution $(P,\sigma,\Delta)$ is safe if $P$ is the set of agents running the protocol; there is a correct $\sigma$-supermajority in $P$, and every $\sigma$-supermajority in $P$ includes a majority of correct agents, hence the intersection of any two $\sigma$-supermajorities contains a correct agent. A constitution is live if $f<\frac{1}{3}n$ and after GST any message sent among correct participants arrives within $\Delta$.
Constitution safety and liveness attest to the validity of the assumption of the constitution regarding the participants and the network connecting them. Note that $\Delta$ bounds the message delay among correct (i.e., participating and online) agents after GST, not the time until every agent comes online. The dual-mode design means the protocol is quiescent when there are no pending transactions, so agents need not be continuously available. When a transaction arrives, only a correct $\sigma$-supermajority must respond within $\Delta$; agents that are temporarily offline are accounted for by the choice of $\sigma$.
@@
Constitutional Consensus problem definition. A constitutional consensus protocol may have multiple instances. Each instance commences with an initial constitution $(P,\sigma,\Delta)$. As in standard consensus, participants spontaneously issue transactions (these are the inputs to the protocol) and every agent outputs a sequence of transactions.
@@
In addition, Constitutional Consensus assumes that participants engage in decision processes as specified by the constitution, which may result in decisions to amend the prevailing constitution. As an example, Section 5 shows how known democratic decision processes can be applied in this new context. Note that the parameters $P$ and $\sigma$ are shared: they govern both the consensus protocol and the democratic decision processes themselves. The interface between the Constitutional Consensus protocol and the constitutional democratic decision processes is as follows: (i) a democratic process may produce constitutional amendment decisions, which become known to the participants once made; (ii) participants can verify that a constitutional amendment decision is valid according to the prevailing constitution, as specified in Section 5. A valid constitutional amendment decision defines a new constitution $(P^{\prime},\sigma^{\prime},\Delta^{\prime})$. An initial constitution is the new constitution of a valid amendment of a constitution with an empty population.
@@
Constitutional amendment decisions are provided to the consensus protocol as inputs by participants that know of them. Thus, the protocol’s output sequence includes both transactions and valid constitutional amendment decisions, as follows:
@@
The constitutional output-sequences produced by correct agents executing a constitutional consensus protocol instance must satisfy the following safety and liveness conditions:
@@
A constitutional consensus protocol is correct if, in any run of it in which all valid constitutional amendments are safe, there exists a constitutional sequence $T=t_{1},t_{2},\dots$ for which the following properties hold:
Consistency:
@@
Validity:
If a correct agent $p$ outputs $t\in T$, then $p$ is a participant in $t$’s epoch.
Epoch Liveness:
For a correct participant $p$ in an epoch $d$ in $T$, $p$’s output eventually includes all the transactions in $d$’s epoch sequence. Furthermore, if $d$ is the last epoch in $T$, then every transaction input by a correct participant in epoch $d$ is included in $T$.
Constitution Amendment Liveness: